ive beans have a mass of 2.1 g. How many beans are in 0.454 kg of beans?

Answer by mananth(16946)

5 beans weigh 2.1 g
..
0.454 kg ?
1kg = 1000 grams
0.454kg = 1000 * 0.454 = 454 grams
..
2.1 gram = 5 beans
454 grams = 454 * 5/2.1
= 1081 beans ( rounded off)